Suppose that Jada Corp. stock is expected to pay a dividend of $0.75 at the end of the year and that the dividend is expected to grow at a rate of 7%. The company’s current beta is 1.9, the current risk-free rate is 2.5% and the market risk premium is 8%. What is the intrinsic value of this stock?

$7.01

None of these

$4.24

$10.71

$7.50

Gertrude Corp bonds have a 13% annual coupon, 7 years left until maturity, and they are currently selling for $1250. What is the yield to maturity of Gertrude’s bonds?

8.00%

15.24%

None of these

8.17%

14.79%
Answer #1

1)

Dividend, D0 = $ 0.75

Growth rate in dividends(g) = 7%

Cost of equity(k) = Risk free rate + Beta * market risk premium =  2.5% + 1.9 * 8% = 17.7 %

Intrinsic Value of stock = D0*(1+g)/(k-g) = 0.75 * (1+ 7%)/(17.7 % - 7%) = 7.5 %

2)

Par value of the bond(FV) = $ 1250

Price of the bond (PV) = 1000

Annual coupon for the bond (PMT) = 13% * 1000 = 130

Number of years to maturity (nper) = 7 years

YTM of the bond = =rate(7,130,-1000, 1250 | RATE(nper, pmt, pv, [fv], [type], [guess]) = 15.24%